OpenStudy (anonymous):

What percent of 105 is 84? i got the answer 5.95238095 but it was wrong can some one help me

OpenStudy (anonymous):

To get the solution, we are looking for, we need to point out what we know. 1. We assume, that the number 105 is 100% - because it's the output value of the task. 2. We assume, that x is the value we are looking for. 3. If 100% equals 105, so we can write it down as 100%=105. 4. We know, that x% equals 84 of the output value, so we can write it down as x%=84. 5. Now we have two simple equations: 1) 100%=105 2) x%=84 where left sides of both of them have the same units, and both right sides have the same units, so we can do something like that: 100%/x%=105/84 6. Now we just have to solve the simple equation, and we will get the solution we are looking for. 7. Solution for 84 is what percent of 105 100%/x%=105/84 (100/x)*x=(105/84)*x - we multiply both sides of the equation by x 100=1.25*x - we divide both sides of the equation by (1.25) to get x 100/1.25=x 80=x x=80 now we have: 84 is 80% of 105

OpenStudy (anonymous):

Let the percent be \(x\)..

OpenStudy (anonymous):

wow that helps so much

OpenStudy (anonymous):

\[x \% of 105 = 84 \\ \frac{x}{100} \times 105 = 84 \\ x = \frac{84 \times 100}{105}\]

OpenStudy (anonymous):

can you help me on this one too I did it and i got 761.71875 32% of 243.75 is what? @eta @rafrasan

OpenStudy (anonymous):

\[\implies \frac{32}{100} \times 243.75 = ??\]

OpenStudy (anonymous):

You check your answer again please..

OpenStudy (anonymous):

i got the same thing

OpenStudy (anonymous):

But I did not..